OpenStudy (anonymous):

simplify 4 radical 20?

OpenStudy (anonymous):

Firstly make the factors of 20. Do it...

OpenStudy (anonymous):

@waterineyes 1, 20, 2, 10, 4, 5 ?

OpenStudy (anonymous):

No dear... See, I give you one example then you will do that for 20. See, I make the factors of 30.. I first divide it by 2, I get 15 as quotient.. So, 30 = 2*15 Now I divide 15 by 3, I get 5 as quotient.. So, 30 = 2*3*5 Now, 5 is not divisible by any other number.. So, factors of 30 = 2*3*5 Now can you make the factors of 20???

OpenStudy (anonymous):

@waterineyes 2 x 2 x 5

OpenStudy (anonymous):

Yes dear you got it very much true.. Now this is in radical.. So, out of two same factors only one come out of radical.. I will give you the example.. 40 = 2*2*2*5 Now, out of 2 two's one will go out of the radical..